The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor-hours. At the beginning of the year, the company provided the following estimates: Department Molding Painting Direct labor-hours 31,500 59,000 Machine-hours 88,000 30,000 Fixed manufacturing overhead cost $ 281,600 $ 566,400 Variable manufacturing overhead per machine-hour $ 2.40 - Variable manufacturing overhead per direct labor-hour - $ 4.40 Job 205 was started on August 1 and completed on August 10. The company's cost records show the following information concerning the job: Department Molding Painting Direct labor-hours 74 126 Machine-hours 310 66 Direct materials $ 938 $ 1,220 Direct labor cost $ 730 $ 960 Required: 1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department. 2. Compute the total overhead cost applied to Job 205. 3-a. What would be the total manufacturing cost recorded for Job 205? 3-b. If the job contained 23 units, what would be the unit product cost?
Answer #1

1) Overhead rate :

Molding department = (281600/88000)+2.40 = 5.60 per machine hour

Painting department (566400/59000)+4.40 = 14 per labor hour

2) Total overhead cost applied = (310*5.6+126*14) = 3500

3a) Total manufacturing cost

Job 205
Direct material (938+1220) 2158
Direct labor (730+960) 1690
Overhead 3500
Total manufacturing cost 7348

3b) Unit product cost = 7348/23 = 319.48 or 319